### v1

Abstract
Despite their proficiency in math tasks, the mechanisms underlying LLMs’ mathematical reasoning abilities remain a subject of debate. Recent studies suggest that chain-of-thought (CoT) prompts can bolster mathematical reasoning by encouraging LLMs to employ human-like logical reasoning (System 2), enabling them to excel on the Cognitive Reflection Test (CRT). To assess whether LLMs genuinely possess System 2-like logical reasoning, we introduced targeted modifications to CRT problems. Our findings reveal that, despite the use of CoT prompts, mainstream LLMs, including the latest o1-preview model, continue to exhibit a significant error rate. Further analysis indicates that they predominantly rely on System 1-like intuitive reasoning and pattern matching derived from training data, rather than demonstrating mastery of mathematical thinking. This discovery challenges the prevailing notion that LLMs possess genuine logical reasoning abilities and that CoT can enhance them. Consequently, this work may temper overly optimistic projections regarding LLMs’ advancement toward artificial general intelligence.
keywords:
Larger Language Model, Chain-of-Thought, Cognitive Reflection Test
1 Introduction
Large language models (LLMs) are hailed as the dawn of Artificial General Intelligence (AGI)[1]. LLMs such as ChatGPT[2], GPT-4[1], Claude[3], Gemini[4], GLM[5], and o1-preview[6] have garnered significant attention from both academia and industry, have demonstrated tremendous potential in various fields such as education[7], healthcare[8, 9], coding[10], and social governance[11]. One of the key reasons for this is the ‘emergence phenomenon’[12], enabled through massive training data and vast parameter counts: LLMs exhibit unexpected competence in tasks they were not specifically trained for.
Using mathematical skills as an illustration, not all LLMs have received specific training for solving complex mathematical problems during their development. Nonetheless, most of them have shown remarkable abilities to tackle these issues. Moreover, by using the CoT method, the ability of LLMs to solve mathematical problems can be further enhanced [13, 14, 15, 16, 17]. However, due to the interpretability challenges posed by large-scale neural networks, there is no scientific consensus yet on the origins of the mathematical capabilities of LLMs. Meanwhile, recent research has questioned the effectiveness of the CoT method, suggesting that CoT provides strong performance benefits primarily on tasks involving math or logic, with much smaller gains on other types of tasks[18].
Specifically, Thilo Hagendorff et al. [19] have demonstrated through experiments that the CoT method can effectively help LLMs handle the pitfalls in Cognitive Reflection Test (CRT) problems[20]. CRT problems are some well-crafted math or logic problems that human testers often get wrong due to intuitive thinking (System 1[21, 22, 23, 24]). By using the CoT method, LLMs are prompted to rely more on human-like logical reasoning (System 2[21, 22, 23, 24]), enhancing their accuracy in solving these problems. Advanced models like GPT-4 have even achieved higher accuracy than humans in these CRT tasks. However, the authors also raised speculative questions in their work, suggesting, “It is possible that some models encountered enough examples in their training to solve them ‘from memory’”[19]. Our study aims to investigate this speculation through empirical research.
We repeated and improved upon the experiment conducted by the previous study [19], making targeted (yet as simple as possible) modifications to the original CRT problems. If LLMs genuinely possess the intrinsic capability to comprehend mathematical logic, as is widely hypothesized, their accuracy in responding to the modified queries should not experience a marked decrement. However, our experiments revealed a distinctly opposing result: even when employing the CoT approach, prominent LLMs, including the latest o1 model, continued to manifest a considerable error rate for the modified problems. Further analysis indicates that LLMs may have not developed a logic reasoning proficiency akin to System 2, nor have they acquired comprehensive mathematical cognitive abilities, including principles of mathematical operations and algorithm selection. They predominantly resort to a methodology reminiscent of System 1, which involves matching and producing responses based on the similarity between user inquiries and training data during the process of text generation. This investigation may serve to temper the overly optimistic anticipations regarding the effectiveness of CoT and the competencies of LLMs in approximating AGI.
2 Experiment I: Changing the numbers in a problem without altering its description and principle
We first replicated the experiment on the CRT3 dataset from Thilo Hagendorff’s study[19], which comprises 50 mathematical test problems. We introduced three types of modifications to the test problems, as delineated in Table 1. Type A represents the original problems in the CRT3 test, characterized as typical exponential growth problems, and incorporates three numbers in the problem statement. Type B modifies two of these numbers, while type C modifies all three. Type D replaces two numbers with letters, transforming the arithmetic problem into an algebraic one. To prevent LLMs from being unsure how to handle algebraic symbols, we appended a suffix for guidance: “X and Y are both numbers, you can use them to represent the final answer.” The modifications in Types B, C, and D do not alter the description of the original problem (Type A). Thus, the underlying mathematical principle of the original problem remains unchanged. To uncover the problem-solving strategies of LLMs, we appended explicit instructions (’Please think step by step’) to the end of each of the four types of problems, encouraging the use of the CoT method.
Table 1: Problem Types and Modifications
Type Modification Example
A Original Problem In a city, a virus is spreading, causing the total number of infected individuals to double each day. If it takes 6 days for the entire city’s population to be infected, how many days would it require for half of the people to become infected? Please think step by step.
B Two Numbers Modified In a city, a virus is spreading, causing the total number of infected individuals to triple each day. If it takes 6 days for the entire city’s population to be infected, how many days would it require for 1/27 of the people to become infected? Please think step by step.
C All Three Numbers Modified In a city, a virus is spreading, causing the total number of infected individuals to triple every three days. If it takes 36 days for the entire city’s population to be infected, how many days would it require for 1/27 of the people to become infected? Please think step by step.
D Algebraic Transformation In a city, a virus is spreading, causing the total number of infected individuals to triple every 3X days. If it takes 6Y days for the entire city’s population to be infected, how many days would it require for 1/27 of the people to become infected? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.
Figure 1: Accuracy of five mainstream LLMs in answering four types of problems.
We evaluated five mainstream open-source and commercial models, manually analyzing and recording their accuracy in responding to the problems. A response was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices Table 2 and Table 3 ). To ensure the reliability of our experimental results, we conducted three replicates of the experiments and computed the average for each data set. The results are shown in Figure 1. In the case of Type A problems, the highest accuracy reached 100.00%, the lowest was 54.00%, and the average accuracy was 86.80%. For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.00%, the lowest 25.33%, and the average 68.53%($\delta=18.27\%;\chi^{2}(1)=23.00;P<0.001$). For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.00%, the lowest at 10.67%, and the average 53.06%($\delta=15.47\%;\chi^{2}(1)=11.91;P<0.001$). The accuracy dropped sharply once again when the arithmetic problems were transformed into algebraic ones (Type D). The best-performing generic model, GPT-4, achieved an accuracy of only 29.33%, whereas ChatGPT 3.5 scarcely provided correct responses to any problems. The other models only got around ten problems right, reducing the average accuracy down to 20.93%($\delta=32.13\%;\chi^{2}(1)=54.00;P<0.001$).
Result Analysis of Experiment I: The modifications in types B, C, and D only altered the specific numbers in the original problems without changing the problems’ mathematical principles and computational rules. For a computer program, merely changing the input parameters without modifying the computation process should not lead to a decrease in accuracy. However, the performance of LLMs differs significantly from that of computer programs. We analyzed all incorrect responses provided by the five LLMs for each type of problem and found that in type B, errors arising from mathematical operations (e.g., 16/2=4) constituted 6.78%, while those due to errors in calculation steps (e.g., omitting a step or altering the original calculation method) accounted for 83.47%, and errors caused by both factors accounted for 9.75%. In type C, errors solely due to mathematical operations accounted for 5.11%, errors due to calculation steps accounted for 85.51%, and errors caused by both factors accounted for 9.38%. In type D, errors due to mathematical operations accounted for 2.19%, errors due to calculation steps accounted for 95.45%, and errors caused by both factors accounted for 2.36%.
The above experimental results indicate that when choosing problem-solving strategies, these LLMs rely not only on the mathematical principles described in textual information of the problems but are also significantly influenced by the numerical values provided in the problem settings. This is obviously different from the human logical reasoning system (system 2).
3 Experiment II: Modifying the principle of a problem while maintaining similarity in its description
In this experiment, we substantially altered the fundamental principles of the mathematical problems, endeavoring to preserve descriptions that closely mirrored the original versions. Subsequently, we investigated whether LLMs persisted in utilizing their prior problem-solving strategies or adapted their methods to the modified problem contexts throughout the problem-solving process. We conducted three distinct sub-experiments corresponding to the CRT1, CRT2, and CRT3 datasets.
3.1 For CRT1: Transforming Additively Separable Problems into Non-Additively Separable Problems
For instance, Figure 2 illustrates the crucial distinction between an original CRT1 problem and its modified version. In the original problem, the total cost is simply the sum of the prices of two items. However, in the modified version, reaching the two items from the starting point through a shared segment must be deducted from the overall distance calculation. We use the string similarity calculation function within the difflib module of Python to quantify the degree of similarity in textual expression between the modified problem and the original one, resulting in an average similarity of 75.91%. Despite the similar wording in the problem statements, the underlying mathematical principles are fundamentally distinct. To reveal how LLM approaches these problems, we included a CoT prompt for both the original and the modified problems.
Figure 2: An example of an original CRT1 problem and its modified problem
We evaluated five mainstream open-source and commercial LLMs and analyzed their accuracy in answering both the original problems and the modified problems. An answer was deemed correct only if both the reasoning steps and the conclusion were accurate. For each problem, we replicated the experiment three times, averaging the results to ensure robustness. Figure 3 illustrates our findings. For the original CRT1 problems, four LLMs achieved a perfect accuracy of 100%, with GLM-4 closely following at 95.33%, yielding an overall accuracy of 99.07%. However, when confronted with the modified problems, the accuracy significantly declined. GLM-4 recorded the highest accuracy rate, albeit a mere 5.33%. Claude3 and Gemini-1.5 failed to provide correct responses to any of the modified problems. On average, the five LLMs attained only 1.47% accuracy($\delta=97.60\%;\chi^{2}(1)=472.41;P<0.001$). By analyzing incorrect answers, we found that 95.30% of errors were attributed to the application of the original problem’s solution method to the modified problem, whereas 4.70% resulted from incorrect application of a new method or computational errors.
Figure 3: Accuracy of the five LLMs when answering the original and the modified CRT1 problems.
3.2 For CRT2: Transforming Problems Related to the number of Individuals into Problems Independent of Individual Count
As illustrated in Figure 4, we take the first problem in the original CRT2 dataset as an example. In the original problem, all workers work together, and the total amount of product is related to the number of workers. In the modified problem, working together is changed to walking together, therefore, the total distance walked is independent of the number of people. We calculated the average similarity using the same function as in the previous sub-experiment, yielding 83.67%. Although the descriptions of the original problem and the modified one are similar, the underlying mathematical principles are entirely different. To uncover the problem-solving strategies of LLMs, we added CoT prompts to both the original and the modified problems.
Figure 4: An example of an original CRT2 problem and its modified problem
Figure 5: Accuracy of the five LLMs when answering the original and the modified CRT2 problems.
As illustrated in Figure 5, the experimental findings demonstrate a significant decrease in the accuracy of various models. Claude3’s accuracy plummeted from an initial level of 62.00% to 22.00% ($\delta=40.00\%;\chi^{2}(1)=14.82;P<0.001$). Similarly, GPT-4 experienced a dramatic reduction from 94.00% to 33.33%($\delta=60.67\%;\chi^{2}(1)=37.20;P<0.001$). Gemini-1.5 experienced the most substantial decline, dropping from 94.67% to 19.33%($\delta=75.34\%;\chi^{2}(1)=54.85;P<0.001$). GLM-4 also encountered a considerable decrease, with its accuracy falling from 65.33% to 28.00%($\delta=37.33\%;\chi^{2}(1)=12.54;P<0.001$). ChatGPT 3.5 exhibited a smaller change, declining from 50.67% to 34.67%($\delta=16.00\%;\chi^{2}(1)=2.00;P=0.157$).
Upon scrutinizing all erroneous answers to the modified problems, it was observed that 52.81% of errors employed the original CRT2 methodology to address the modified problems. This finding suggests that the LLMs perceive the modified problems as akin to the originals due to their descriptive similarity, consequently relying on the original solution strategy.
3.3 For CRT3: Transforming Exponential Growth Problems into Linear Growth Problems
Figure 6 presents the first problem in the CRT3 dataset as an illustrative example. In the original problem, the number of viruses doubles each day, resulting in exponential growth of the total number. In contrast, the modified problem features a constant daily increase in the number of viruses, which leads to linear growth. We calculated the average similarity using the same function, yielding a result of 88.89%. To further elucidate the problem-solving strategies of the LLM, we added a CoT prompt for both the original and the modified problem.
Figure 6: An example of an original CRT3 problem and its modified version
The experimental results in Figure 7 demonstrate that Gemini-1.5 achieved the highest accuracy of 100% for the original CRT3 problems. In contrast, ChatGPT 3.5 recorded the lowest accuracy at 54%, with an average accuracy of 86.80%. Following the modification of the problems, the average accuracy decreased significantly to 12.53% ($\delta=74.27\%;\chi^{2}(1)=272.83;P<0.001$). The accuracy rate of Gemini-1.5 decreased to 20.00%($\delta=80.00\%;\chi^{2}(1)=63.38;P<0.001$), while Claude3’s accuracy was merely 1.33% ($\delta=92.00\%;\chi^{2}(1)=81.23;P<0.001$).
The principal reason for the decline in the average accuracy is that these models continued to employ the problem-solving approach of the original problem when confronted with the modified problem. The proportions of Claude3, ChatGPT 3.5, GPT-4, Gemini-1.5, and GLM-4 that identified the modified problem as one of exponential growth were 81.69%, 77.90%, 85.88%, 72.49%, and 71.49%, respectively, although it should have been identified as a linear growth problem. Among these, the proportion of problems that were solved entirely using the original method—yielding the same answer as the original problem—constituted 37.85%, 31.83%, 82.70%, 55.73%, and 52.17% of all incorrect answers, with an average of 52.06%. This result shows that more than half of the errors were due to using the original method.
Figure 7: Accuracy of the five LLMs when answering the original and the modified CRT3 problems.
Result Analysis of Experiment II: The modifications to CRT1, CRT2, and CRT3 altered the mathematical principles underlying these problems with a effort to preserve the similarity of their statements. The experimental findings reveal that LLMs continue to rely on their inherent problem-solving methodologies, even when provided with CoT prompts as guidance. This observation further substantiates our hypothesis that generative LLMs operate similarly to a word chain game; they can address some mathematical issues probably because their training datasets include analogous problems. Consequently, LLMs tend to adopt problem-solving strategies based on the similarity of problem statements rather than a genuine understanding of the mathematical principles involved.
4 Experiment III: Replicating Experiment I & II on the Latest o1 Model
Given OpenAI’s latest release of the o1-preview model, which focuses on enhancing logical reasoning capabilities [6], we have repeated the above experiments with o1. The results are shown in Figure 8.
In terms of replicating Experiment I, changing the numbers in the problem statements did not significantly affect the accuracy of o1’s answers, which aligns with expectations. This observation suggests that (1) o1 has enhanced numerical computation capabilities compared to previous LLMs, and (2) the specific number given in the problem statement does not influence the method o1 chooses to solve the problem. This might imply that o1 has potentially incorporated prompts like “list the equations before solving” into its built-in thought process. Nevertheless, this speculation cannot be officially confirmed, as OpenAI has not released technical details regarding o1’s improved reasoning capabilities.
However, when replicating Experiment II, the average accuracy rate of o1 was only 10%, which was even lower than that of most previous LLMs. After making minimal textual modifications while altering the mathematical principles underlying the original CRT1, CRT2, and CRT3 problems, o1 persisted in selecting problem-solving approaches based on the mathematical principles corresponding to the original problems. The errors stemming from this persistence constituted 100%, 68.42%, and 75.86% of the incorrect responses to the three types of CRT problems, respectively. This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets, cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems. The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes (such as those involving word chains), leading to their thought patterns being deeply ingrained.
Figure 8: Accuracy of the o1 model in Experiment I and Experiment II.
5 Conclusion
This paper draws on the classic CRT problems from human psychology to conduct an empirical study on the “emergence” of mathematical capabilities in mainstream LLMs. By constructing forward experiments (Experiment 1) and reverse experiments (Experiment 2), the conclusions were starkly different from mainstream views. Specifically:
(1) LLMs tend to match problem-solving strategies based on textual similarity rather than truly understanding the principles of mathematical problems.
(2) Even with the introduction of the CoT reasoning or improving reasoning ability through specialized training (e.g., o1), such methods cannot entirely adjust the LLMs’ inherent but incorrect problem-solving approaches.
This paper focuses on mathematical ability, which is the foundation of all scientific research, as its central point of study. It provides a counterexample to the overexpectation of the “emergence phenomenon” of LLM capabilities and the CoT method, both of which are currently receiving much attention.
Declarations
To enhance academic communication and promote related research, we will open-source our dataset and experimental data at https://osf.io/74yj2/.
In order to facilitate testing, we chose to use the unified API provided by Huiyan Company to call the official APIs of various LLMs.
Appendix A Appendix
Table 2: Examples of Determining Incorrect Answers for Four Types of Problems
Problem Type Incorrect Answer Example
A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt? Please think step by step.”, ”answer”: ”If the colony doubles in size every day, then on the first day it covers half of the yogurt. On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered. Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”}
B {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Let’s break down the problem step by step: 1. The colony triples daily: This means that if the initial size of the colony is 1 unit, it will be 3 units the next day, 9 units the following day, 27 units the day after, and so on. 2. It takes 4 days for the colony to completely cover the yogurt: This means that the colony reaches its maximum size on the 4th day, which is 27 units. 3. We need to find out how many days it takes for the colony to cover 1/27 of the yogurt: Since the colony triples in size daily, on the third day, the colony would be 9 units, which is 1/3 of the complete coverage. So, it means that on the third day the yogurt is 1/3 covered. Therefore, it takes 3 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 13:22:38”, ”Model”: ”gpt-3.5-turbo-0125”}
C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days. If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”First, let’s figure out how many times the colony triples in size in the 27 days. Since the colony triples every three days, it will triple in size 27/3 = 9 times. Now, we know that the colony starts at 1/27 of the yogurt and triples in size 9 times. So, to find out how many days it takes for the colony to cover 1/27 of the yogurt, we need to divide 27 by 9, which equals 3. Therefore, it takes 3 days for the colony to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 15:55:15”, ”Model”: ”gpt-3.5-turbo-0125”}
D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days. If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.”, ”answer”: ”To find out how many days it takes for the patch to cover 1/27 of the yogurt, we can first find the number of times the colony triples in size to cover the entire yogurt. If the colony triples in size every 3X days, then in 4Y days, the colony would have tripled in size 4Y / 3X times. To find out how many days it takes for the colony to cover 1/27 of the yogurt, we can use the concept of exponential growth. If the colony triples every 3X days, then after n periods of 3X days, the size of the colony would be $(3^{n})$ times the original size. So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(}4Y/3X)$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt. $3^{(}4Y/3X)=3^{(}-3)$ Since the bases are the same, we can set the exponents equal to each other. 4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”}
Table 3: Examples of Determining Correct Answers for Four Types of Problems
Problem Type Correct Answer Example
A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt? Please think step by step.”, ”answer”: ”Let’s think about the growth of the bacteria colony step by step: 1. On the first day, the colony is so small that it is only a fraction of the yogurt’s surface. Let’s say it is 1/2 of its final size (since it doubles each day). 2. On the second day, the colony doubles in size. Now it is 1/2 * 2 = 1 times its final size, or 100% of its final size. 3. Since it takes 4 days for the colony to completely cover the yogurt, we know that on the third day it must be at 100% of its final size because it doubles each day. 4. Now, if on the third day the colony is at full size, on the second day it must be at half the size of the third day (since it doubles each day). Therefore, the colony covers half of the yogurt on the third day from the starting point. So, it takes 3 days for the patch to cover half of the yogurt.”, ”time”: ”2024-08-23 09:23:15”, ”Model”: ”glm-4”}
B {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Let’s think about the growth of the bacteria colony step by step. We know that the colony triples in size daily. If it takes 4 days to completely cover the yogurt, then on the fourth day, the colony is at its maximum size. Now, let’s consider the size of the colony on each day leading up to the fourth day: - On the first day, the colony is at 1/3 of its final size because it will triple each day.- On the second day, the colony is at 1/3 of the size it was on the fourth day, but since it tripled from the first day, it’s now at 1/3 * 3 = 1/9 of the final size.- On the third day, it triples again, so it’s at 1/9 * 3 = 1/3 of the final size.- On the fourth day, it triples one last time to reach its full size, which means it’s at 1/3 * 3 = 1, or 100% of the final size. We’re interested in when the colony covers 1/27 of the yogurt. Since the colony triples each day, we can work backward from the fourth day: - If on the fourth day the colony is at 100% (or 1), then on the third day it was at 1/3 of that size, which is 1/3 * 1 = 1/3.- If on the third day it was at 1/3, then on the second day it was at 1/3 of that size, which is 1/3 * 1/3 = 1/9.- Continuing this pattern, on the first day, the colony would be at 1/3 of the size it was on the second day, which is 1/3 * 1/9 = 1/27. Therefore, it takes 1 day for the colony to reach 1/27 of the yogurt’s surface. The patch covers 1/27 of the yogurt on the very first day of growth.”, ”time”: ”2024-08-13 13:23:05”, ”Model”: ”glm-4”}
Problem Type Correct Answer Example
C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days. If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Let’s think step by step through the problem: 1. We know that the size of the colony triples every three days. This means that if we represent the size of the colony at any time as ( C ), then after three days, the size of the colony will be ( 3C ). 2. Since the colony triples in size every three days, we can say that the size of the colony increases by a factor of ( 3 ) each time interval of three days passes. So, if it takes 27 days for the colony to completely cover the yogurt, then we can divide 27 by 3 to find out how many times the colony triples in size during this period: ( $frac\{27text\{days\}\}\{3text\{days/triple\}\}=9text\{triples\}$ ) This tells us that the colony triples in size 9 times over the course of 27 days. 3. Now, if the colony triples 9 times to cover the entire yogurt, we need to find out how many times it would triple to cover ( $frac\{1\}\{27\}$ ) of the yogurt. Since ( $frac\{1\}\{27\}$ ) is ( $3^{\{}-3\}$ ), or ( $frac\{1\}\{3^{3}\}$ ), we can infer that we need to triple the colony size three times fewer than the full coverage:($9text\{triples\}-3text\{triples\}=6text\{triples\}$) This means the colony needs to triple in size 6 times to cover ( $frac\{1\}\{27\}$) of the yogurt. 4. Finally, we need to calculate how many days it takes for the colony to triple in size 6 times. Since each triple takes three days, we can multiply the number of triples by three: ( 6 text{ triples} times 3 text{ days/triple} = 18 text{ days} ) So, it takes 18 days for the patch to cover ( $frac\{1\}\{27\}$) of the yogurt.”, ”time”: ”2024-08-13 15:56:43”, ”Model”: ”glm-4”}
D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days. If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.”, ”answer”: ”Let’s break down the problem step by step: 1. The colony triples every 3X days. This means that if we let ( T ) represent the number of days it takes for one patch to form, then after ( T ) days, we have 1 patch, after ( T + 3X ) days, we have 3 patches (because it tripled), after ( T + 6X ) days, we have 9 patches (3 times 3), and so on. 2. After 4Y days, the colony completely covers the yogurt. This means that after 4Y days, we must have ( $3^{n}$ ) patches, where ( n ) is an integer, and ( $3^{n}$ ) is equal to the total number of patches required to cover the yogurt. 3. To find out how many patches are needed to cover 1/27 of the yogurt, we can use the fact that ( 1/27 ) is ( $3^{-3}$ ) (since ( $273^{3}=$ 27 ) and ( 1/27 ) is ( 27 ) to the power of ( -1 )). 4. Now we need to find the time it takes to reach this number of patches. Since the colony triples every ( 3X ) days, we need to find out how many times we can triple the colony to get to ( $3^{-3}$ ) times the total number of patches needed to cover the yogurt. 5. If it takes 4Y days to completely cover the yogurt, then it takes ( 4Y - 3X ) days to have ( $3^{n-1}$ ) patches, one step before complete coverage. Similarly, it takes ( 4Y - 6X ) days to have ( $3^{n-2}$ ) patches, and so on. 6. To find the time it takes to reach ( $3^{-3}$ ) of the total patches, we need to go back 3 steps from complete coverage. This means: ( 4Y - 9X ) days to have ( $3^{n-3}$ ) patches (since each step is ( 3X ) days). 7. But we know that ( $3^{n-3}$ ) is equal to ( $3^{-3}$ ) because ( $3^{n-3}$ ) is ( $3^{n}$ ) divided by ( $3^{3}$ ), and ( $3^{n}$) is the number of patches needed for full coverage. 8. Therefore, ( 4Y - 9X ) is the number of days it takes to cover ( 1/27 ) of the yogurt. Final Answer: It takes ( 4Y - 9X ) days for the patch to cover ( 1/27 ) of the yogurt.”, ”time”: ”2024-08-14 20:53:14”, ”Model”: ”glm-4”}

### v2

@@
 We repeated and improved upon the experiment conducted by the previous study [19], making targeted (yet as simple as possible) modifications to the original CRT problems.
-If LLMs genuinely possess the intrinsic capability to comprehend mathematical logic, as is widely hypothesized, their accuracy in responding to the modified queries should not experience a marked decrement.
+If LLMs genuinely possess the intrinsic capability to comprehend mathematical logic, as is widely hypothesized, their accuracy in responding to the modified problems should not experience a marked decrement.
 However, our experiments revealed a distinctly opposing result: even when employing the CoT approach, prominent LLMs, including the latest o1 model, continued to manifest a considerable error rate for the modified problems.
@@
 Thus, the underlying mathematical principle of the original problem remains unchanged.
-To uncover the problem-solving strategies of LLMs, we appended explicit instructions (’Please think step by step’) to the end of each of the four types of problems, encouraging the use of the CoT method.
+To uncover the problem-solving strategies of LLMs, we appended explicit instructions (’Please think step by step.’) to the end of each of the four types of problems, encouraging the use of the CoT method.
 Table 1: Problem Types and Modifications
@@
 We evaluated five mainstream open-source and commercial models, manually analyzing and recording their accuracy in responding to the problems.
-A response was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices Table 2 and Table 3 ).
+A response was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices A ).
 To ensure the reliability of our experimental results, we conducted three replicates of the experiments and computed the average for each data set.
@@
 For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.00%, the lowest 25.33%, and the average 68.53%($\delta=18.27\%;\chi^{2}(1)=23.00;P<0.001$).
-For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.00%, the lowest at 10.67%, and the average 53.06%($\delta=15.47\%;\chi^{2}(1)=11.91;P<0.001$).
+For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.00%, the lowest at 10.66%, and the average 53.06%($\delta=15.47\%;\chi^{2}(1)=11.91;P<0.001$).
 The accuracy dropped sharply once again when the arithmetic problems were transformed into algebraic ones (Type D).
@@
 This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets, cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems.
-The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes (such as those involving word chains), leading to their thought patterns being deeply ingrained.
+The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes (such as auto-regressive next-token predictors), leading to their thought patterns being deeply ingrained.
 Figure 8: Accuracy of the o1 model in Experiment I and Experiment II.
@@
 In order to facilitate testing, we chose to use the unified API provided by Huiyan Company to call the official APIs of various LLMs.
-Appendix A Appendix
+Appendix A Examples
 Table 2: Examples of Determining Incorrect Answers for Four Types of Problems
-Problem Type Incorrect Answer Example
+Problem Type Incorrect Answer Example Error Type
 A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily.
@@
 On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered.
-Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”}
+Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
 B {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples daily.
@@
 So, it means that on the third day the yogurt is 1/3 covered.
-Therefore, it takes 3 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 13:22:38”, ”Model”: ”gpt-3.5-turbo-0125”}
+Therefore, it takes 3 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 13:22:38”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
 C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days.
@@
 So, to find out how many days it takes for the colony to cover 1/27 of the yogurt, we need to divide 27 by 9, which equals 3.
-Therefore, it takes 3 days for the colony to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 15:55:15”, ”Model”: ”gpt-3.5-turbo-0125”}
+Therefore, it takes 3 days for the colony to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 15:55:15”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
 D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days.
@@
 So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(4Y/3X)}$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt.
-$3^{(}4Y/3X)=3^{(}-3)$ Since the bases are the same, we can set the exponents equal to each other.
-4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”}
+$3^{(4Y/3X)}=3^{-3}$ Since the bases are the same, we can set the exponents equal to each other.
+4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps and mathematical operations
 Table 3: Examples of Determining Correct Answers for Four Types of Problems
@@
 Problem Type Correct Answer Example
-C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days.
-If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt?
-Please think step by step.”, ”answer”: ”Let’s think step by step through the problem: 1.
-We know that the size of the colony triples every three days.
-This means that if we represent the size of the colony at any time as ( C ), then after three days, the size of the colony will be ( 3C ).
-2.
-Since the colony triples in size every three days, we can say that the size of the colony increases by a factor of ( 3 ) each time interval of three days passes.
-So, if it takes 27 days for the colony to completely cover the yogurt, then we can divide 27 by 3 to find out how many times the colony triples in size during this period: ( $frac\{27text\{days\}\}\{3text\{days/triple\}\}=9text\{triples\}$ ) This tells us that the colony triples in size 9 times over the course of 27 days.
-3.
-Now, if the colony triples 9 times to cover the entire yogurt, we need to find out how many times it would triple to cover ( $frac\{1\}\{27\}$ ) of the yogurt.
-Since ( $frac\{1\}\{27\}$ ) is ( $3^{\{}-3\}$ ), or ( $frac\{1\}\{3^{3}\}$ ), we can infer that we need to triple the colony size three times fewer than the full coverage:($9text\{triples\}-3text\{triples\}=6text\{triples\}$) This means the colony needs to triple in size 6 times to cover ( $frac\{1\}\{27\}$) of the yogurt.
-4.
-Finally, we need to calculate how many days it takes for the colony to triple in size 6 times.
-Since each triple takes three days, we can multiply the number of triples by three: ( 6 text{ triples} times 3 text{ days/triple} = 18 text{ days} ) So, it takes 18 days for the patch to cover ( $frac\{1\}\{27\}$) of the yogurt.”, ”time”: ”2024-08-13 15:56:43”, ”Model”: ”glm-4”}
+[…]
 D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days.

### v4

@@
 Abstract
-Despite their proficiency in math tasks, the mechanisms underlying LLMs’ mathematical reasoning abilities remain a subject of debate.
-Recent studies suggest that chain-of-thought (CoT) prompts can bolster mathematical reasoning by encouraging LLMs to employ human-like logical reasoning (System 2), enabling them to excel on the Cognitive Reflection Test (CRT).
-To assess whether LLMs genuinely possess System 2-like logical reasoning, we introduced targeted modifications to CRT problems.
-Our findings reveal that, despite the use of CoT prompts, mainstream LLMs, including the latest o1-preview model, continue to exhibit a significant error rate.
-Further analysis indicates that they predominantly rely on System 1-like intuitive reasoning and pattern matching derived from training data, rather than demonstrating mastery of mathematical thinking.
-This discovery challenges the prevailing notion that LLMs possess genuine logical reasoning abilities and that CoT can enhance them.
-Consequently, this work may temper overly optimistic projections regarding LLMs’ advancement toward artificial general intelligence.
+The cognitive mechanism by which Large Language Models (LLMs) solve mathematical problems remains a widely debated and unresolved issue.
+Currently, there is little interpretable experimental evidence that connects LLMs’ problem-solving with human cognitive psychology.
+To determine if LLMs possess human-like mathematical reasoning, we modified the problems used in the human Cognitive Reflection Test (CRT).
+Our results show that, even with the use of Chains of Thought (CoT) prompts, mainstream LLMs, including the latest o1 model (noted for its reasoning capabilities), have a high error rate when solving these modified CRT problems.
+Specifically, the average accuracy rate dropped by up to 50% compared to the original questions.
+Further analysis of LLMs’ incorrect answers suggests that they primarily rely on pattern matching from their training data, which aligns more with human intuition (System 1 thinking) rather than with human-like reasoning (System 2 thinking).
+This finding challenges the belief that LLMs have genuine mathematical reasoning abilities comparable to humans.
+As a result, this work may adjust overly optimistic views on LLMs’ progress towards artificial general intelligence.
 keywords:
@@
 1 Introduction
-Large language models (LLMs) are hailed as the dawn of Artificial General Intelligence (AGI)[1].
-LLMs such as ChatGPT[2], GPT-4[1], Claude[3], Gemini[4], GLM[5], and o1-preview[6] have garnered significant attention from both academia and industry, have demonstrated tremendous potential in various fields such as education[7], healthcare[8, 9], coding[10], and social governance[11].
-One of the key reasons for this is the ‘emergence phenomenon’[12], enabled through massive training data and vast parameter counts: LLMs exhibit unexpected competence in tasks they were not specifically trained for.
-Using mathematical skills as an illustration, not all LLMs have received specific training for solving complex mathematical problems during their development.
-Nonetheless, most of them have shown remarkable abilities to tackle these issues.
-Moreover, by using the CoT method, the ability of LLMs to solve mathematical problems can be further enhanced [13, 14, 15, 16, 17].
-However, due to the interpretability challenges posed by large-scale neural networks, there is no scientific consensus yet on the origins of the mathematical capabilities of LLMs.
-Meanwhile, recent research has questioned the effectiveness of the CoT method, suggesting that CoT provides strong performance benefits primarily on tasks involving math or logic, with much smaller gains on other types of tasks[18].
-Specifically, Thilo Hagendorff et al.
-[19] have demonstrated through experiments that the CoT method can effectively help LLMs handle the pitfalls in Cognitive Reflection Test (CRT) problems[20].
+Large language models (LLMs) are considered the dawn of Artificial General Intelligence (AGI) [1].
+Models such as ChatGPT[2], GPT-4[1], Claude[4], Gemini[3], GLM[5], and o1-preview[23] have garnered considerable attention from both academia and industry.
+These models exhibit significant potential across various fields, including education[6], healthcare[7, 8], coding[9], and social governance[10].
+This is partly attributed to the ‘emergence phenomenon’ [12], which allows LLMs, due to their large training datasets and numerous parameters, to perform tasks they were not specifically trained for.
+Using mathematical skills as an example, most LLMs have demonstrated remarkable abilities to tackle these problems.
+By leveraging the Chain of Thought (CoT) method, the capacity of LLMs to solve mathematical problems can be further augmented[11, 20, 21, 19, 22].
+However, due to the interpretability challenges posed by large-scale neural networks, there is still no scientific consensus on the origins and mechanisms underlying the mathematical capabilities of LLMs.
+Conducting psychological measurement experiments on LLMs can help enhance the interpretability of research into LLMs’ thinking.
+Thilo Hagendorff et al.[13] have demonstrated through experiments that the CoT method can effectively help LLMs handle the pitfalls in Cognitive Reflection Test (CRT) problems[14].
 CRT problems are some well-crafted math or logic problems that human testers often get wrong due to intuitive thinking (System 1[15, 16, 18, 17]).
@@
 Advanced models like GPT-4 have even achieved higher accuracy than humans in these CRT tasks.
-However, the authors also raised speculative questions in their work, suggesting, “It is possible that some models encountered enough examples in their training to solve them ‘from memory’”[19].
-Our study aims to investigate this speculation through empirical research.
+However, the authors also raised speculative questions in their work, suggesting, “It is possible that some models encountered enough examples in their training to solve them ‘from memory’”.
 We repeated and improved upon the experiment conducted by the previous study[13], making targeted (yet as simple as possible) modifications to the original CRT problems.
@@
 However, our experiments revealed a distinctly opposing result: even when employing the CoT approach, prominent LLMs, including the latest o1 model, continued to manifest a considerable error rate for the modified problems.
-Further analysis indicates that LLMs may have not developed a logic reasoning proficiency akin to System 2, nor have they acquired comprehensive mathematical cognitive abilities, including principles of mathematical operations and algorithm selection.
+Further analysis of the incorrect answers indicates that LLMs may have not developed a logic reasoning proficiency akin to System 2, nor have they acquired comprehensive mathematical cognitive abilities.
 They predominantly resort to a methodology reminiscent of System 1, which involves matching and producing responses based on the similarity between user inquiries and training data during the process of text generation.
@@
 2 Experiment I: Changing the Numbers in A Problem Without Altering Its Description and Principle
-We first replicated the experiment on the CRT3 dataset from Thilo Hagendorff’s study[19], which comprises 50 mathematical test problems.
+We first replicated the experiment on the CRT3 dataset from Thilo Hagendorff’s study[13], which comprises 50 mathematical problems.
 We introduced three types of modifications to the test problems, as delineated in Table 1.
@@
 Thus, the underlying mathematical principle of the original problem remains unchanged.
-To uncover the problem-solving strategies of LLMs, we appended explicit instructions (’Please think step by step.’) to the end of each of the four types of problems, encouraging the use of the CoT method.
+To uncover the problem-solving strategies of LLMs, we appended explicit instructions (‘Please think step by step.’) to the end of each of the four types of problems, encouraging the use of the CoT method.
 Table 1: Problem Types and Modifications
@@
 If it takes 6Y days for the entire city’s population to be infected, how many days would it require for 1/27 of the people to become infected?
-X and Y are both numbers, you can use them to represent the final answer.
+X and Y are both numbers, and you can use them to represent the final answer.
 Please think step by step.
 Figure 1: Accuracy of five mainstream LLMs in answering four types of problems.
-We evaluated five mainstream open-source and commercial models, manually analyzing and recording their accuracy in responding to the problems.
-A response was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices A ).
+We evaluated five mainstream open-source and commercial LLMs by manually analyzing and recording their answers to the problems, ensuring the precision of our result judgments.
+An answer was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices A ).
 To ensure the reliability of our experimental results, we conducted three replicates of the experiments and computed the average for each data set.
-The results are shown in Figure 1.
-In the case of Type A problems, the highest accuracy reached 100.00%, the lowest was 54.00%, and the average accuracy was 86.80%.
-For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.00%, the lowest 25.33%, and the average 68.53%($\delta=18.27\%;\chi^{2}(1)=23.00;P<0.001$).
-For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.00%, the lowest at 10.66%, and the average 53.06%($\delta=15.47\%;\chi^{2}(1)=11.91;P<0.001$).
+The results are shown in Fig.
+1.
+In the case of Type A problems, the highest accuracy reached 100%, the lowest was 54.0%, and the average accuracy was 86.8%.
+For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.0%, the lowest 25.3%, and the average 68.5%($\delta=18.3\%;\chi^{2}(1)=23.00;P<0.001$).
+For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.0%, the lowest at 10.7%, and the average 53.1%($\delta=15.4\%;\chi^{2}(1)=11.91;P<0.001$).
 The accuracy dropped sharply once again when the arithmetic problems were transformed into algebraic ones (Type D).
-The best-performing generic model, GPT-4, achieved an accuracy of only 29.33%, whereas ChatGPT 3.5 scarcely provided correct responses to any problems.
-The other models only got around ten problems right, reducing the average accuracy down to 20.93%($\delta=32.13\%;\chi^{2}(1)=54.00;P<0.001$).
+The best-performing model, GPT-4, achieved an accuracy of only 29.3%, whereas ChatGPT 3.5 scarcely provided correct answers to any problems.
+The average accuracy was reduced to 20.9%($\delta=32.2\%;\chi^{2}(1)=54.00;P<0.001$).
 Result Analysis of Experiment I: The modifications in types B, C, and D only altered the specific numbers in the original problems without changing the problems’ mathematical principles and computational rules.
-For a computer program, merely changing the input parameters without modifying the computation process should not lead to a decrease in accuracy.
-However, the performance of LLMs differs significantly from that of computer programs.
-We analyzed all incorrect responses provided by the five LLMs for each type of problem and found that in type B, errors arising from mathematical operations (e.g., 16/2=4) constituted 6.78%, while those due to errors in calculation steps (e.g., omitting a step or altering the original calculation method) accounted for 83.47%, and errors caused by both factors accounted for 9.75%.
-In type C, errors solely due to mathematical operations accounted for 5.11%, errors due to calculation steps accounted for 85.51%, and errors caused by both factors accounted for 9.38%.
-In type D, errors due to mathematical operations accounted for 2.19%, errors due to calculation steps accounted for 95.45%, and errors caused by both factors accounted for 2.36%.
+For a human who possesses mathematical logical thinking and is capable of solving problems through formulating equations or programming, merely changing the input numbers without modifying the mathematic principles should not lead to a decrease in accuracy.
+However, the performance of LLMs differs significantly from that of a rational human.
+As shown in Fig.
+2, we analyzed all incorrect answers provided by the five LLMs for each type of problem and found that in type B, errors arising from incorrect solution steps (e.g., omitting a step or altering the original calculation method) constituted 93.2%, while those only due to mathematical calculation (e.g., 16/2=4) accounted for 6.8%.
+In type C, errors due to incorrect solution steps accounted for 94.9%, errors solely due to mathematical calculation accounted for 5.1%.
+In type D, errors due to incorrect solution steps accounted for 97.8%, and errors caused solely by mathematical calculation accounted for 2.2%.
+Figure 2: Analysis of the proportion of incorrect answers in Experiment I
 The above experimental results indicate that when choosing problem-solving strategies, these LLMs rely not only on the mathematical principles described in textual information of the problems but are also significantly influenced by the numerical values provided in the problem settings.
@@
 3 Experiment II: Modifying the Principle of A Problem While Maintaining Similarity in Its Description
+To further substantiate the inferences drawn from Experiment I, we designed a reverse experiment.
 In this experiment, we substantially altered the fundamental principles of the mathematical problems, endeavoring to preserve descriptions that closely mirrored the original versions.
@@
 3.1 For CRT1: Transforming Additively Separable Problems into Non-Additively Separable Problems
-For instance, Figure 2 illustrates the crucial distinction between an original CRT1 problem and its modified version.
+For instance, Fig.
+3 illustrates the crucial distinction between an original CRT1 problem and its modified version.
 In the original problem, the total cost is simply the sum of the prices of two items.
@@
 For each problem, we replicated the experiment three times, averaging the results to ensure robustness.
-Figure 3 illustrates our findings.
-For the original CRT1 problems, four LLMs achieved a perfect accuracy of 100%, with GLM-4 closely following at 95.33%, yielding an overall accuracy of 99.07%.
+Fig.
+4A illustrates our findings.
+For the original CRT1 problems, four LLMs achieved a perfect accuracy of 100%, with GLM-4 closely following at 95.3%, yielding an overall accuracy of 99.1%.
 However, when confronted with the modified problems, the accuracy significantly declined.
-GLM-4 recorded the highest accuracy rate, albeit a mere 5.33%.
-Claude3 and Gemini-1.5 failed to provide correct responses to any of the modified problems.
-On average, the five LLMs attained only 1.47% accuracy($\delta=97.60\%;\chi^{2}(1)=472.41;P<0.001$).
-By analyzing incorrect answers, we found that 95.30% of errors were attributed to the application of the original problem’s solution method to the modified problem, whereas 4.70% resulted from incorrect application of a new method or computational errors.
-Figure 3: Accuracy of the five LLMs when answering the original and the modified CRT1 problems.
+GLM-4 recorded the highest accuracy rate, albeit a mere 5.3%.
+Claude3 and Gemini-1.5 failed to provide correct answers to any of the modified problems.
+On average, the five LLMs attained only 1.5% accuracy($\delta=97.6\%;\chi^{2}(1)=472.41;P<0.001$).
+Figure 4: Accuracy of the LLMs when answering the CRT1 problems (A), and the proportion of incorrect answer types in modified CRT1 problems (B).
+By analyzing incorrect answers, as shown in Fig.
+4B, we found that 95.9% of errors were attributed to the application of the original problem’s solution method to the modified problem (Maintain the original solution steps), whereas 4.1% resulted from new solution steps but with errors.
 3.2 For CRT2: Transforming Problems Related to the Number of Individuals into Problems Independent of Individual Count
-As illustrated in Figure 4, we take the first problem in the original CRT2 dataset as an example.
+As illustrated in Fig.
+5, we take the first problem in the original CRT2 dataset as an example.
 In the original problem, all workers work together, and the total amount of product is related to the number of workers.
-In the modified problem, working together is changed to walking together, therefore, the total distance walked is independent of the number of people.
-We calculated the average similarity using the same function as in the previous sub-experiment, yielding 83.67%.
+In the modified problem, working together is changed to walking together.
+Therefore, the total distance walked is independent of the number of people.
+We calculated the average similarity using the same function as in the previous sub-experiment, yielding 82.33%.
 Although the descriptions of the original problem and the modified one are similar, the underlying mathematical principles are entirely different.
@@
 Figure 5: An example of an original CRT2 problem and its modified problem
-Figure 5: Accuracy of the five LLMs when answering the original and the modified CRT2 problems.
-As illustrated in Figure 5, the experimental findings demonstrate a significant decrease in the accuracy of various models.
-Claude3’s accuracy plummeted from an initial level of 62.00% to 22.00% ($\delta=40.00\%;\chi^{2}(1)=14.82;P<0.001$).
-Similarly, GPT-4 experienced a dramatic reduction from 94.00% to 33.33%($\delta=60.67\%;\chi^{2}(1)=37.20;P<0.001$).
-Gemini-1.5 experienced the most substantial decline, dropping from 94.67% to 19.33%($\delta=75.34\%;\chi^{2}(1)=54.85;P<0.001$).
-GLM-4 also encountered a considerable decrease, with its accuracy falling from 65.33% to 28.00%($\delta=37.33\%;\chi^{2}(1)=12.54;P<0.001$).
-ChatGPT 3.5 exhibited a smaller change, declining from 50.67% to 34.67%($\delta=16.00\%;\chi^{2}(1)=2.00;P=0.157$).
-Upon scrutinizing all erroneous answers to the modified problems, it was observed that 52.81% of errors employed the original CRT2 methodology to address the modified problems.
-This finding suggests that the LLMs perceive the modified problems as akin to the originals due to their descriptive similarity, consequently relying on the original solution strategy.
+As illustrated in Fig.
+6A, the experimental findings demonstrate a significant decrease in the accuracy of various models.
+Claude3’s accuracy plummeted from an initial level of 62.0% to 22.0% ($\delta=40.0\%;\chi^{2}(1)=14.82;P<0.001$).
+Similarly, GPT-4 experienced a dramatic reduction from 94.0% to 33.3%($\delta=60.7\%;\chi^{2}(1)=37.20;P<0.001$).
+Gemini-1.5 experienced the most substantial decline, dropping from 94.7% to 19.3%($\delta=75.4\%;\chi^{2}(1)=54.85;P<0.001$).
+GLM-4 also encountered a considerable decrease, with its accuracy falling from 65.3% to 28.0%($\delta=37.33\%;\chi^{2}(1)=12.54;P<0.001$).
+ChatGPT 3.5 exhibited a smaller change, declining from 50.7% to 34.7%($\delta=16.0\%;\chi^{2}(1)=2.00;P=0.157$).
+Figure 6: Accuracy of the LLMs when answering the CRT2 problems (A), and the proportion of incorrect answer types in modified CRT2 problems (B).
+Upon scrutinizing all erroneous answers to the modified problems, as shown in Fig.
+6B, it was observed that 52.8% of errors employed the original CRT2 methodology to address the modified problems (Maintain the original solution steps).
+This indicates that, due to the similarity in wording between the modified problem and the original problem, LLMs overlooked the differences in their mathematical principles and consequently chose the same solution steps as those for the original problem, which was incorrect.
 3.3 For CRT3: Transforming Exponential Growth Problems into Linear Growth Problems
-Figure 6 presents the first problem in the CRT3 dataset as an illustrative example.
+Fig.
+7 presents the first problem in the CRT3 dataset as an illustrative example.
 In the original problem, the number of viruses doubles each day, resulting in exponential growth of the total number.
@@
 Figure 7: An example of an original CRT3 problem and its modified version
-The experimental results in Figure 7 demonstrate that Gemini-1.5 achieved the highest accuracy of 100% for the original CRT3 problems.
-In contrast, ChatGPT 3.5 recorded the lowest accuracy at 54%, with an average accuracy of 86.80%.
-Following the modification of the problems, the average accuracy decreased significantly to 12.53% ($\delta=74.27\%;\chi^{2}(1)=272.83;P<0.001$).
-The accuracy rate of Gemini-1.5 decreased to 20.00%($\delta=80.00\%;\chi^{2}(1)=63.38;P<0.001$), while Claude3’s accuracy was merely 1.33% ($\delta=92.00\%;\chi^{2}(1)=81.23;P<0.001$).
-The principal reason for the decline in the average accuracy is that these models continued to employ the problem-solving approach of the original problem when confronted with the modified problem.
-The proportions of Claude3, ChatGPT 3.5, GPT-4, Gemini-1.5, and GLM-4 that identified the modified problem as one of exponential growth were 81.69%, 77.90%, 85.88%, 72.49%, and 71.49%, respectively, although it should have been identified as a linear growth problem.
-Among these, the proportion of problems that were solved entirely using the original method—yielding the same answer as the original problem—constituted 37.85%, 31.83%, 82.70%, 55.73%, and 52.17% of all incorrect answers, with an average of 52.06%.
-This result shows that more than half of the errors were due to using the original method.
-Figure 7: Accuracy of the five LLMs when answering the original and the modified CRT3 problems.
+The experimental results in Fig.
+8A demonstrate that Gemini-1.5 achieved the highest accuracy of 100% for the original CRT3 problems.
+In contrast, ChatGPT 3.5 recorded the lowest accuracy at 54.0%, with an average accuracy of 86.8%.
+Following the modification of the problems, the average accuracy decreased significantly to 12.5% ($\delta=64.3\%;\chi^{2}(1)=272.83;P<0.001$).
+The accuracy rate of Gemini-1.5 decreased to 20.0%($\delta=80.0\%;\chi^{2}(1)=63.38;P<0.001$), while Claude3’s accuracy was merely 1.3% ($\delta=92.0\%;\chi^{2}(1)=81.23;P<0.001$).
+Figure 8: Accuracy of the LLMs when answering the CRT3 problems (A), and the proportion of incorrect answer types in modified CRT3 problems (B).
+The main reason for the decline in the average accuracy is that these models continued to employ the problem-solving approach of the original problem when confronted with the modified problem.
+The proportions of Claude3, ChatGPT 3.5, GPT-4, Gemini-1.5, and GLM-4 that identified the modified problem as one of exponential growth were 81.7%, 77.8%, 85.7%, 72.5%, and 71.5%, respectively, although it should have been identified as a linear growth problem.
+Among these, the proportion of problems that were solved entirely using the original method—yielding the same answer as the original problem—constituted 37.8%, 31.8%, 82.7%, 55.7%, 52.2% and the average is 51.4% (as depicted in Fig.
+8B).
+This result indicates that more than half (51.4%) of the errors were due to maintaining the original solution steps.
 Result Analysis of Experiment II: The modifications to CRT1, CRT2, and CRT3 altered the mathematical principles underlying these problems with a effort to preserve the similarity of their statements.
 The experimental findings reveal that LLMs continue to rely on their inherent problem-solving methodologies, even when provided with CoT prompts as guidance.
-This observation further substantiates our hypothesis that generative LLMs operate similarly to a word chain game; they can address some mathematical issues probably because their training datasets include analogous problems.
-Consequently, LLMs tend to adopt problem-solving strategies based on the similarity of problem statements rather than a genuine understanding of the mathematical principles involved.
+This observation further supports our hypothesis that LLMs are capable of addressing mathematical problems, likely due to the inclusion of analogous problems in their training datasets.
+Consequently, LLMs select solutions based on the superficial similarity of problem descriptions, rather than demonstrating a deep understanding of the underlying mathematical principles.
 4 Experiment III: Replicating Experiment I & II on the Latest o1 Model
-Given OpenAI’s latest release of the o1-preview model, which focuses on enhancing logical reasoning capabilities [6], we have repeated the above experiments with o1. The results are shown in Figure 8.
-In terms of replicating Experiment I, changing the numbers in the problem statements did not significantly affect the accuracy of o1’s answers, which aligns with expectations.
-This observation suggests that (1) o1 has enhanced numerical computation capabilities compared to previous LLMs, and (2) the specific number given in the problem statement does not influence the method o1 chooses to solve the problem.
+Given OpenAI’s latest release of the o1-preview model, which focuses on enhancing logical reasoning capabilities[23], we have repeated the above experiments with o1. The results are shown in Fig. 9.
+In terms of replicating Experiment I (Fig.
+9A), changing the numbers in the problem statements did not significantly affect the accuracy of o1’s answers.
+This observation suggests that the specific number given in the problem statement does not affect the method that o1 selects for solving the problem.
 This might imply that o1 has potentially incorporated prompts like “list the equations before solving” into its built-in thought process.
 Nevertheless, this speculation cannot be officially confirmed, as OpenAI has not released technical details regarding o1’s improved reasoning capabilities.
-However, when replicating Experiment II, the average accuracy rate of o1 was only 10%, which was even lower than that of most previous LLMs.
-After making minimal textual modifications while altering the mathematical principles underlying the original CRT1, CRT2, and CRT3 problems, o1 persisted in selecting problem-solving approaches based on the mathematical principles corresponding to the original problems.
-The errors stemming from this persistence constituted 100%, 68.42%, and 75.86% of the incorrect responses to the three types of CRT problems, respectively.
-This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets, cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems.
-The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes (such as auto-regressive next-token predictors), leading to their thought patterns being deeply ingrained.
-Figure 8: Accuracy of the o1 model in Experiment I and Experiment II.
+However, when replicating Experiment II (Fig.
+9B), the average accuracy rate of o1 was only 10.0%.
+After the mathematical principles of the CRT problem are altered while modifying the textual description as little as possible, o1 persisted in selecting problem-solving approaches based on the mathematical principles corresponding to the original problems.
+The errors stemming from this persistence constituted 100%, 68.4%, and 75.7% of the incorrect answers to the three types of CRT problems, respectively.
+As shown in the Fig.
+9C, on average, 82.5% of the incorrect answers were due to treating the modified problem as the original one and providing the same answer as for the original problem.
+This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets (e.g.
+the o1 model), cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems.
+The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes, such as the adoption of auto-regressive next-token predictors, resulting in their thought patterns being deeply ingrained, akin to human intuition (System 1) rather than logical reasoning (System 2).
+Figure 9: Average Accuracy of the o1 model in Experiment I (A) and Experiment II (B), and the average proportion of incorrect answers in Experiment II (C).
+5 Related Works
+Some previous empirical research has questioned the reasoning abilities of LLMs[26, 25, 27, 24], particularly in solving mathematical problems.
+In contrast, the contributions of this paper are more pronounced in the following aspects:
+Most of the previous related studies were published before the release of the o1 model (e.g.[26, 27, 24]), therefore did not assess this model, which is renowned for its reasoning abilities.
+The latest experimental evidence provided in this paper demonstrates that similar conclusions remain valid for the newly released o1 model.
+Although some previous studies have found that some LLMs exhibit overfitting and memorization effects during the reasoning process, leading to a decline in performance on modified datasets, the degree of decline was generally not significant (e.g.[27] , less than 13%), and not all LLMs experienced a decrease.
+Consequently, these studies were cautious in drawing conclusions, given that even for human testers, a moderate decline is inevitable.
+However, in this paper, the accuracy of all LLMs drops by more than 50% on average, providing more convincing experimental evidence.
+We also analyze the incorrect answers generated by LLMs, demonstrating that the reasons for their decline in accuracy are significantly different from those of humans.
+Some previous studies have also mentioned human cognitive science to compare and analyze the thinking patterns of LLMs[26, 25].
+However, their datasets used for testing mathematical reasoning ability are purely mathematical and unrelated to cognitive science itself.
+In contrast, our dataset itself is a mathematical test dataset from cognitive science, originally used to study the roles of human reasoning (System 2) and intuition (System 1).
+Therefore, this paper offers greater interpretability and inspiration from a psychological perspective.
 6 Conclusion
 This paper draws on the classic CRT problems from human psychology to conduct an empirical study on the “emergence” of mathematical capabilities in mainstream LLMs.
-By constructing forward experiments (Experiment 1) and reverse experiments (Experiment 2), the conclusions were starkly different from mainstream views.
+It aims to test whether LLMs possess genuine mathematical reasoning capabilities and to provide more interpretable explanations for the causes from the perspective of human psychology.
+By constructing forward experiments (Experiment I) and reverse experiments (Experiment II), we obtained conclusions that are starkly different from mainstream views.
 Specifically:
-(1) LLMs tend to match problem-solving strategies based on textual similarity rather than truly understanding the principles of mathematical problems.
-(2) Even with the introduction of the CoT reasoning or improving reasoning ability through specialized training (e.g., o1), such methods cannot entirely adjust the LLMs’ inherent but incorrect problem-solving approaches.
-This paper focuses on mathematical ability, which is the foundation of all scientific research, as its central point of study.
-It provides a counterexample to the overexpectation of the “emergence phenomenon” of LLM capabilities and the CoT method, both of which are currently receiving much attention.
+  • LLMs tend to match problem-solving strategies based on textual similarity rather than truly understanding the underlying principles of mathematical problems.
+This process is more akin to human intuition (System 1) rather than logical reasoning (System 2).
+  • Even with the introduction of CoT or specialized training to enhance reasoning abilities (e.g., o1), such methods cannot fundamentally alter the problem-solving thinking patterns of LLMs to endow them with human-like logical reasoning abilities (System 2).
+This is because the prevailing paradigm for training and fine-tuning LLMs (i.e., generating the next token with the highest probability) is, at its core, similar to human intuition (System 1).
+This study conducts an empirical analysis of the “emergence” of LLMs’ mathematical reasoning abilities from a psychological perspective.
+We hope that this study can reduce overblown expectations of LLMs’ capabilities and stimulate more empirical research to objectively evaluate the current limitations of LLMs’ abilities.
 Declarations
@@
 Appendix A Examples
-Table 2: Examples of Determining Incorrect Answers for Four Types of Problems
-Problem Type Incorrect Answer Example Error Type
-A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily.
-If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt?
-Please think step by step.”, ”answer”: ”If the colony doubles in size every day, then on the first day it covers half of the yogurt.
-On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered.
-Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
-B {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples daily.
-If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt?
-Please think step by step.”, ”answer”: ”Let’s break down the problem step by step: 1.
-The colony triples daily: This means that if the initial size of the colony is 1 unit, it will be 3 units the next day, 9 units the following day, 27 units the day after, and so on.
-2.
-It takes 4 days for the colony to completely cover the yogurt: This means that the colony reaches its maximum size on the 4th day, which is 27 units.
-3.
-We need to find out how many days it takes for the colony to cover 1/27 of the yogurt: Since the colony triples in size daily, on the third day, the colony would be 9 units, which is 1/3 of the complete coverage.
-So, it means that on the third day the yogurt is 1/3 covered.
-Therefore, it takes 3 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 13:22:38”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
-C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days.
-If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt?
-Please think step by step.”, ”answer”: ”First, let’s figure out how many times the colony triples in size in the 27 days.
-Since the colony triples every three days, it will triple in size 27/3 = 9 times.
-Now, we know that the colony starts at 1/27 of the yogurt and triples in size 9 times.
-So, to find out how many days it takes for the colony to cover 1/27 of the yogurt, we need to divide 27 by 9, which equals 3.
-Therefore, it takes 3 days for the colony to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 15:55:15”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
-D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days.
-If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt?
-X and Y are both numbers, you can use them to represent the final answer.
-Please think step by step.”, ”answer”: ”To find out how many days it takes for the patch to cover 1/27 of the yogurt, we can first find the number of times the colony triples in size to cover the entire yogurt.
-If the colony triples in size every 3X days, then in 4Y days, the colony would have tripled in size 4Y / 3X times.
-To find out how many days it takes for the colony to cover 1/27 of the yogurt, we can use the concept of exponential growth.
-If the colony triples every 3X days, then after n periods of 3X days, the size of the colony would be $(3^{n})$ times the original size.
-So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(4Y/3X)}$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt.
-$3^{(4Y/3X)}=3^{-3}$ Since the bases are the same, we can set the exponents equal to each other.
-4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps and mathematical operations
-Table 3: Examples of Determining Correct Answers for Four Types of Problems
+Table 2: Examples of Determining Correct Answers for Four Types of Problems - I
 Problem Type Correct Answer Example
@@
 The patch covers 1/27 of the yogurt on the very first day of growth.”, ”time”: ”2024-08-13 13:23:05”, ”Model”: ”glm-4”}
+Table 3: Examples of Determining Correct Answers for Four Types of Problems - II
 Problem Type Correct Answer Example
@@
 Final Answer: It takes ( 4Y - 9X ) days for the patch to cover ( 1/27 ) of the yogurt.”, ”time”: ”2024-08-14 20:53:14”, ”Model”: ”glm-4”}
+Table 4: Examples of Determining Incorrect Answers for Four Types of Problems - I
+Problem Type Incorrect Answer Example Error Type
+A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily.
+If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt?
+Please think step by step.”, ”answer”: ”If the colony doubles in size every day, then on the first day it covers half of the yogurt.
+On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered.
+Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”} incorrect solution steps
+B {”index”: 33, ”query”: ”An iceberg is forming and its surface triples every year.
+If the iceberg takes 10 years to grow to 27 square miles, how many years are required for it to grow to 1 square miles?
+Please think step by step.”, ”answer”: ”First, let’s find out how much the surface area of the iceberg increases each year.
+If the surface triples every year, then we can represent the surface area of the iceberg over the years as follows: 1 year: x 2 years: 3x 3 years: 3*3x = 9x 4 years: 3*9x = 27x From this pattern, we see that the surface area of the iceberg increases by a factor of 3 each year, so we can represent the surface area after n years as $3^{n}*x$.
+We are given that after 10 years, the iceberg has a surface area of 27 square miles, so we can set up the equation: $27=3^{1}0*x$ Solving for x: $x=27/3^{1}0$ Now that we know the value of x, we can use it to find out how many years it takes for the iceberg to grow to 1 square mile: $1=3^{n}*(27/3^{10})$ Solving for n: $3^{n}=1/(27/3^{10})$, $273^{n}=1*3^{10}/$ 27, $273^{n}=3^{10}/$ 27, $n=log3(3^{10}/27)$; $n\approx 5.07$ So, it takes approximately 5.07 years for the iceberg to grow to 1 square mile.”, ”time”: ”2024-08-20 18:08:33”, ”Model”: ”gpt-3.5-turbo-0125”} mathematical calculation errors
+C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days.
+If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt?
+Please think step by step.”, ”answer”: ”Here’s how to solve this problem step-by-step: **1.
+Work Backwards:** * Since the colony triples in size every three days, it means it was 1/3 the size three days before it was full.
+* If it takes 27 days to cover the yogurt, it means it was 1/3 the size on day 24 (27 - 3).
+* It was 1/3 the size again on day 21 (24 - 3).
+* And again on day 18 (21 - 3).
+* And again on day 15 (18 - 3).
+* And again on day 12 (15 - 3).
+* And again on day 9 (12 - 3).
+* And again on day 6 (9 - 3).
+* And again on day 3 (6 - 3).
+**2.
+Identify the 1/27 Point:** * Since the colony triples in size, to be 1/27 of the full size, it means the colony was 1/3 * 1/3 * 1/3 = 1/27 the size three days before it was 1/3 the size.
+**3.
+Find the Answer:** * We found that the colony was 1/3 the size on day 3.
+* Three days before that (day 0) the colony was 1/27 the size.
+**Therefore, it takes 0 days for the colony to cover 1/27 of the yogurt.
+It was already that size at the start.** ”, ”time”: ”2024-08-13 15:56:05”, ”Model”: ”gemini-1.5-flash-latest”} incorrect solution steps
+Table 5: Examples of Determining Incorrect Answers for Four Types of Problems - II
+Problem Type Incorrect Answer Example Error Type
+D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days.
+If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt?
+X and Y are both numbers, you can use them to represent the final answer.
+Please think step by step.”, ”answer”: ”To find out how many days it takes for the patch to cover 1/27 of the yogurt, we can first find the number of times the colony triples in size to cover the entire yogurt.
+If the colony triples in size every 3X days, then in 4Y days, the colony would have tripled in size 4Y / 3X times.
+To find out how many days it takes for the colony to cover 1/27 of the yogurt, we can use the concept of exponential growth.
+If the colony triples every 3X days, then after n periods of 3X days, the size of the colony would be $(3^{n})$ times the original size.
+So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(4Y/3X)}$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt.
+$3^{(4Y/3X)}=3^{-3}$ Since the bases are the same, we can set the exponents equal to each other.
+4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”} incorrect solution steps